# A Certifiable Machine Learning-Based Pipeline to Predict Fatigue Life of Aircraft Structures

### v1

@@ 2.3 Dataset and split description
As previously discussed, missions are divided into operational segments such as taxi, climb, cruise, descent, and approach, being each mission segment described in the dataset by 12 scalar flight parameters collected into a vector $\mathbf{P}\in\mathbb{R}^{12}$. These flight parameters induce certain stresses at each PSE (FEM output), represented by 4-dimensional stress vectors $\mathbf{S}\in\mathbb{R}^{4}$, which represent the local stress at the component for different loading conditions. Details about P and S provided in subsection A.1. We leverage this data to address phase I from subsection 2.1.
@@
Data split – Instead of performing a random split on all the $38\times 7\times 4=1064$ $\{\text{PSE},\text{Mission},k_{t}\}$ combinations- which could result in data leakage and in some PSEs being absent from the test set -the following split was performed. For each PSE, a set of five out of the seven missions have been randomly selected for the training set (for the four $k_{t}$ values, treated as numerical variables), one mission for the validation set and another mission for the test set. The validation set is used for hyperparameter tuning while the test set is used to assess the overall performance of the pipeline. Details on how this split is applied in each phase are provided in subsection A.2 and subsection B.3.
@@ 2.4 Note on software
ML models were developed using pyLOM [22], a high-performance computing-enabled framework for parallel reduced-order modeling implemented in PyTorch [23]. In addition, we also used standard Python libraries, such as scikit-learn [24] for polynomial regression and data preprocessing, pandas [25] and NumPy [26] for dataset manipulation, and matplotlib [27] for visualization.
@@ 3.1 Phase I: Stress predictions
where $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$ are the polynomial coefficients, obtained by minimizing the residual sum of squares between predictions and ground truth values (ordinary least squares) [24].
@@ 3.2 Phase II: GAG and G&M Damage prediction
In Figure 9 and Figure 9 we show a summary of the relative error (%) statistics of the damage predictions $\mathscr{D}_{\text{GAG}}$ and $\mathscr{D}_{\text{G\&M}}$, for all the points in the test set. As we can see, the mean relative error is moderately large and substantially larger that the median, what suggests a skewed, fat-tailed error distribution. Notice, indeed, that the third quartile is not symmetric with respect to the first quartile, and that the max error is disproportionally large. This suggests that pointwise error statistics are being deformed –and the average performance apparently decreased– due to the presence of outliers. So it is important to ascertain where these outliers are located. Since the models predict fatigue damage –which can lead to instantaneous breakage–, greater damages represent more critical situations in both cases, as fatigue failure will be achieved earlier, resulting in shorter fatigue lives. Therefore, it is in the high-damage region that the MLP models must be more accurate, and therefore we need to certify that outliers are never present in this region.
@@
Figure 8: Relative error statistics for the prediction of $\mathscr{D}_{\text{GAG}}$.
@@
Figure 9: Relative error statistics for the prediction of $\mathscr{D}_{\text{G\&M}}$.
Figure 10: Scatter plot of the relative error of $\mathscr{D}_{\text{GAG}}$ as a function of its ground truth value, for every point in the test set.
Figure 11: Scatter plot of the relative error of $\mathscr{D}_{\text{G\&M}}$ as a function of its ground truth value, for every point in the test set.
@@
where $m$ are the number of samples on the mini batch and $y_{i}$, $\hat{y}_{i}$ are the generic true and predicted damages, respectively. By construction, this loss function already penalizes error for higher damages than for lower damages. To confirm this dependence of the prediction’s error on the actual damage, in Figure 10 and Figure 11 we plot, in log-log scales, the relative error for the prediction of accumulated damages $\mathscr{D}_{\text{GAG}}$ and $\mathscr{D}_{\text{G\&M}}$ as a function of their ground truth values. Each point in these two scatter plots represents a $\{\text{Mission},\text{PSE},k_{t}\}\in\text{Test set}$. Clearly, model’s performance is better in the region of high damages (region of interest), and it is in the region of low damages that the error outliers tend to concentrate (observe that the plots are logarithmic). This bias is even stronger for the GAG damage predictions –which is actually the main source of fatigue damage of the studied aircraft–, showing a statistically significant weak inverse correlation between the accumulated damage and the relative error of the prediction (Spearman’s $\rho=-0.353$ with p-value $6.7\times 10^{-5}$). In contrast, for the accumulated damage caused by gusts and maneuvers, the correlation is very weak and not statistically significant (Spearman’s $\rho=-0.17$ with p-value =0.085=0.085). However, the distribution outliers (e.g., following Tukey fences, those points with relative error larger than $Q3+1.5IQR=17.3\%$) are mainly located in the low-damage range.
@@ 3.3 Phase III: Fatigue life results
Miner’s Rule is finally applied to obtain total fatigue damage after the $n$ flights, and with Equation 3 fatigue life $N$ can easily be calculated. Since the MLP models for $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$ showed better performance for higher fatigue-related damages, errors propagate and thus we expect to get more accurate predictions of fatigue life for shorter to medium fatigue lives, $N\in(10^{3}-10^{6})$- which are the most critical situations- since $N\propto\mathscr{D}^{-1}$.
In Figure 12 we scatter plot the the relative error of each fatigue life prediction $N$ as a function of its ground truth values. It is clear that the model shows a good performance in the region of interest (short to medium fatigue lives). Note that the behavior is similar to that of the inverse of Figure 10. This is expected as the main source of damage for the studied aircraft lies in the GAG cycle.
@@
For completeness, in Figure 13 we also show the mean relative error of fatigue life prediction conditioned on $k_{t}$, mission and PSE, respectively. As observed in 13(a), cases with smaller $k_{t}$ concentrate larger relative error. This is a sensible result: small $k_{t}$ values in general tend to have longer fatigue lives (i.e. the region where the ML model performs worse) because in these cases the local maximum stress is barely amplified by defects or geometry characteristics. Likewise, in 13(c) we observe that some PSEs show poorer fatigue life prediction, but this is simply because these are subjected to lower stresses, leading to higher fatigue lives and thus poorer model’s performance. Incidentally, note that PSE 15 was not included in the plot as its high error was off-scale. Finally, 13(b) shows only slight variability in performance across missions.
@@
Finally, Table 6 depicts a summary of the main relative error statistics for fatigue life predictions and in Table 6, the same statistics are presented but for the region of interest $N\in(10^{3}-10^{6})$. In this latter region, performance is highly improved and predictions become more accurate.
@@ 3.4 Uncertainty quantification: prediction intervals
Since our ML pipeline aims to provide a solution for industrial and safety-critical applications, authorities require not only accuracy, but confidence in its predictions. In fatigue analysis, it us customary to use scatter factors to account for the variability in material fatigue properties and the uncertainties inherent in predicting fatigue life. To that end, following the principles of [6] we now build prediction intervals for fatigue life predictions.
We start by considering the relative error of the predictions of the calibration set in the region of interest, $N\in(10^{3}-10^{6})$. Since the relative error is larger or equal to zero, we construct bootstrapped CI95 confidence interval of 95th percentile of this metric [6]. In Figure 14 we plot the bootstrap average of this percentile, such that the range $0,\overline{P95}]$ in principle covers roughly the relative error of about $95\%$ of the points in the calibration and test set. The analysis showed that, using the validation set described in subsection 2.3 as the calibration set to build a prediction interval for the 95% coverage, we achieve a coverage in the test set of 95.24%, CI95 [88.25, 98.69]. This corresponds to an upper bound of the relative error of 13.50%.
@@ 3.5 A posteriori training/test split analysis
Before assessing the performance of the different ML models, it is a good practice to check whether the training/test data split is adequate, as incorrect splits can be a source of p-hacking –which misleadingly amplifies the performance on the test set– or poor generalization [6]. Now, since in this work the output data of model $f_{1}$ is the input data of model $f_{2}$, we now perform such data split analysis a posteriori, and focus in the input data of MLP models $f_{2}$ and $f_{3}$, as the configuration space is smaller than the one for $f_{1}$ (note that this analysis was performed for all steps):
@@
where $f_{2}=\{f_{2}^{\text{GAG}},f_{2}^{\text{G\&M}}\}$. Following [6], we consider a split to be adequate if (i) finding the same data samples appearing in both the train and test sets has vanishingly small probability, (ii) the feature vector distributions of both sets are similar, and (iii) the test set data is adequately close from the training set.
Criterion (ii) is checked by performing 2-sample hypothesis tests (2-sample Kolmogorov-Smirnov [28], 2-sample Anderson-Darling [29] and 2-sample chi-2 tests) on marginal training and test distributions for each input variable. Criterion (iii) is implemented in the so-called tesselation and proximity method [6], by detecting test set points that are either excessively close (statistically speaking) to the training set (source of p-hacking), or statistically too far (isolated points with poor generalization).
@@
Figure 17: Training vs test marginal distributions for the output variable $N$.
In Figure 15 we plot the marginal training and test set distributions for each numerical input variable, whereas in Figure 16 we plot the ones for the categorical variables. Although variables in Figure 16 are actually numerical, for the sake of this analysis they are treated as categorical as they only take a small set of possible values. Therefore, they cannot be considered continuous and a different test is applied ($\chi^{2}$). Likewise, Figure 17 presents the marginal training and test distributions for the output variable $N$. Visually, the training and test distributions are similar, although small differences exist. To quantify the significance of these, we have run several hypothesis tests, where the null hypothesis H0 is that, for each variable, each pair of populations are drawn from the same distribution. Since each hypothesis test has its own null hypothesis, no correction for multiple testing is required, and we adopt the standard choice for rejecting the null. In Table 7 we present the p-values of these tests, where the null hypothesis cannot be rejected for any variable except for $n$ and $t_{flight}$. Note nonetheless that, as observed in Figure 16, the range values for these two variables is similar in both sets, indicating that at least there are no isolated values in $n$ or $t_{flight}$ in the test set.
@@ 4 Conclusions
Looking ahead, while the current study is limited to wing components due to data availability, the methodology is extensible to other materials and aircraft structures, such as the fuselage and stabilizers. Future work should also focus on broader structural coverage, improved stress modeling, and further integration of uncertainty quantification conditioned on input space [6].
@@
Acknowledgments – The authors acknowledge funding from project TIFON (PLEC2023-010251) funded by MCIN/AEI/10.13039/501100011033, Spain. LL acknowledges partial support from projects MISLAND (PID2020-114324GB-C22), CSxAI (PID2024-157526NB-I00) and MdM Seal of Excellence (CEX2021-001164-M) funded by MICIU/AEI/10.13039/501100011033 and from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL). Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer.
@@ B.5 $f_{2}^{\text{G\&M}}$: input, output variables and model specs
In this case we only have 7 input variables: $k_{t},\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4},n,t_{flight}$. The rationale of each variable is similar to that for $f_{2}^{\text{GAG}}$; $\langle 1g\rangle_{\text{flight}}$ shows the stress level around which the oscillations that cause this damage take place. The magnitude of these oscillations is captured by $\langle{\Delta\text{vgust}}\rangle,\langle{\Delta\text{vman}}\rangle,\langle{\Delta\text{turn}}\rangle$ and $n$, $t_{flight}$ have the same statistical impact. Besides, longer flights typically imply more gusts encountered and maneuvers executed (not necessarily larger) and consequently greater damage. The output variable is the per-flight damage $\overline{\mathscr{D}_{\text{G\&M}}}$, such that the cumulated damage is simply $\mathscr{D}_{\text{G\&M}}=n\cdot\overline{\mathscr{D}_{\text{G\&M}}}$. Once again, variables were normalized with MinMax Scaler and logarithm was applied to the damage. The MLP specs and hyperparameters are specified in Table 14. Hyperparameter tuning was performed manually with the validation set to ensure generalization.

### v2

@@ 2.3 Dataset and split description
As previously discussed, missions are divided into operational segments such as taxi, climb, cruise, descent, and approach, being each mission segment described in the dataset by 12 scalar flight parameters collected into a vector $\mathbf{P}\in\mathbb{R}^{12}$. These flight parameters induce certain stresses at each PSE (FEM output), represented by 4-dimensional stress vectors $\mathbf{S}\in\mathbb{R}^{4}$, which represent the local stress at the component for different loading conditions. Details about P and S provided in A.1. We leverage this data to address phase I from subsection 2.1.
@@
Data split – Instead of performing a random split on all the $38\times 7\times 4=1064$ $\{\text{PSE},\text{Mission},k_{t}\}$ combinations- which could result in data leakage and in some PSEs being absent from the test set -the following split was performed. For each PSE, a set of five out of the seven missions have been randomly selected for the training set (for the four $k_{t}$ values, treated as numerical variables), one mission for the validation set and another mission for the test set. The validation set is used for hyperparameter tuning while the test set is used to assess the overall performance of the pipeline. Details on how this split is applied in each phase are provided in A.2 and B.3.
@@ 2.4 Note on software
ML models were developed using pyLOM [pyLOM], a high-performance computing-enabled framework for parallel reduced-order modeling implemented in PyTorch [paszke2019pytorch]. In addition, we also used standard Python libraries, such as scikit-learn [scikit-learn] for polynomial regression and data preprocessing, pandas [mckinney-proc-scipy-2010] and NumPy [harris2020array] for dataset manipulation, and matplotlib [Hunter:2007] for visualization.
@@ 3.1 Phase I: Stress predictions
where $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$ are the polynomial coefficients, obtained by minimizing the residual sum of squares between predictions and ground truth values (ordinary least squares) [scikit-learn].
@@ 3.2 Phase II: GAG and G&M Damage prediction
In Figure 9 and Figure 9 we show a summary of the relative error (%) statistics of the damage predictions $\mathbb{D}_{\text{GAG}}$ and $\mathbb{D}_{\text{G\&M}}$, for all the points in the test set. As we can see, the mean relative error is moderately large and substantially larger that the median, what suggests a skewed, fat-tailed error distribution. Notice, indeed, that the third quartile is not symmetric with respect to the first quartile, and that the max error is disproportionally large. This suggests that pointwise error statistics are being deformed –and the average performance apparently decreased– due to the presence of outliers. So it is important to ascertain where these outliers are located. Since the models predict fatigue damage –which can lead to instantaneous breakage–, greater damages represent more critical situations in both cases, as fatigue failure will be achieved earlier, resulting in shorter fatigue lives. Therefore, it is in the high-damage region that the MLP models must be more accurate, and therefore we need to certify that outliers are never present in this region.
@@
Figure 8: Relative error statistics for the prediction of $\mathbb{D}_{\text{GAG}}$.
@@
Figure 9: Relative error statistics for the prediction of $\mathbb{D}_{\text{G\&M}}$.
Figure 10: Scatter plot of the relative error of $\mathbb{D}_{\text{GAG}}$ as a function of its ground truth value, for every point in the test set.
Figure 11: Scatter plot of the relative error of $\mathbb{D}_{\text{G\&M}}$ as a function of its ground truth value, for every point in the test set.
@@
where $m$ are the number of samples on the mini batch and $y_{i}$, $\hat{y}_{i}$ are the true and predicted damages, respectively. By construction, this loss function already penalizes error for higher damages than for lower damages. To confirm this dependence of the prediction’s error on the actual damage, in Figure 10 and Figure 11 we plot, in log-log scales, the relative error for the prediction of accumulated damages $\mathbb{D}_{\text{GAG}}$ and $\mathbb{D}_{\text{G\&M}}$ as a function of their ground truth values. Each point in these two scatter plots represents a $\{\text{Mission},\text{PSE},k_{t}\}\in\text{Test set}$. Clearly, model’s performance is better for higher damages (more critical situations), and it is in the region of low damages that the error outliers tend to concentrate (observe that the plots are logarithmic). This bias is even stronger for the GAG damage predictions –which is actually the main source of fatigue damage of the studied aircraft–, showing a statistically significant weak inverse correlation between the accumulated damage and the relative error of the prediction (Spearman’s $\rho=-0.353$ with p-value $6.7\times 10^{-5}$). In contrast, for the accumulated damage caused by gusts and maneuvers, the correlation is very weak and not statistically significant (Spearman’s $\rho=-0.17$ with p-value =0.085=0.085). However, the distribution outliers (e.g., following Tukey fences, those points with relative error larger than $Q3+1.5IQR=17.3\%$) are mainly located in the low-damage range.
@@ 3.3 Phase III: Fatigue life results
Miner’s Rule is finally applied to obtain total fatigue damage after the $n$ flights, and with Eq.3 fatigue life NN can easily be calculated. Since the MLP models for $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$ showed poorer performance for lower fatigue-related damages, errors propagate and thus we expect to get less accurate predictions of fatigue life for higher fatigue lives, since $N\propto\mathbb{D}^{-1}$.
In Figure 12 we scatter plot the relative error of each fatigue life prediction NN as a function of its ground truth values. As observed, errors of the predictions are not uniformly distributed along fatigue life. For lives over $10^{6}$ the model shows poorer performance. This behavior is similar to that of the inverse of Figure 10. This is expected as the main source of damage for the studied aircraft lies in the GAG cycle. This poorer performance for lives over $10^{6}$ flights does not represent a significant limitation for our model, since these very long lives do not represent critical situations and big errors do not pose a safety hazard. In contrast, shorter lives are more critical, as failure will be reached earlier in the aircraft operation.
@@
For completeness, in Figure 13 we also show the mean relative error of fatigue life prediction conditioned on $k_{t}$, mission and PSE, respectively. As observed in 13(a), cases with smaller $k_{t}$ concentrate larger relative error. This is a sensible result: small $k_{t}$ values in general tend to have longer fatigue lives (i.e. the region where the ML model performs worse) because in these cases the local maximum stress is barely amplified by defects or geometry characteristics.
Likewise, in 13(c) we observe that some PSEs show poorer fatigue life prediction, but this is simply because these are subjected to lower stresses, leading to higher fatigue lives and thus poorer model’s performance. Incidentally, note that PSE 15 was not included in the plot as its high error was off-scale. Finally, 13(b) shows only slight variability in performance across missions.
@@
Finally, Table 6 depicts a summary of the main relative error statistics for fatigue life predictions and in Table 6, the same statistics are presented excluding samples with lives exceeding $10^{6}$ flights, since these are non-critical and some of them can highly perturb the error statistics. In this latter region, performance is highly improved and predictions become more accurate.
@@ 3.4 Uncertainty quantification: prediction intervals
Since our ML pipeline aims to provide a solution for industrial and safety-critical applications, authorities require not only accuracy, but confidence in its predictions. In fatigue analysis, it us customary to use scatter factors to account for the variability in material fatigue properties and the uncertainties inherent in predicting fatigue life. To that end, following the principles of [lacasa2025certification] we now build prediction intervals for fatigue life predictions.
We start by considering the relative error of the predictions of the calibration set in the region $N\in(10^{3}-10^{6})$ to avoid impractical situations that can perturb the error statistics. Since the relative error is larger or equal to zero, we construct bootstrapped confidence interval of 95th percentile (CI95) of this metric [lacasa2025certification]. In Figure 14 we plot the bootstrap average of this percentile, such that the range $0,\overline{P95}]$ in principle covers roughly the relative error of about $95\%$ of the points in the calibration and test set. The analysis showed that, using the validation set described in subsection 2.3 as the calibration set to build a prediction interval for the 95% coverage, we achieve a coverage in the test set of 95.24%, CI95 [88.25, 98.69]. This corresponds to an upper bound of the relative error of 13.50%.
@@ 3.5 A posteriori training/test split analysis
Before assessing the performance of the different ML models, it is a good practice to check whether the training/test data split is adequate, as incorrect splits can be a source of p-hacking –which misleadingly amplifies the performance on the test set– or poor generalization [lacasa2025certification]. Now, since in this work the output data of model $f_{1}$ is the input data of model $f_{2}$, we now perform such data split analysis a posteriori, and focus in the input data of MLP models $f_{2}$ and $f_{3}$, as the configuration space is smaller than the one for $f_{1}$ (note that this analysis was performed for all steps):
@@
where $f_{2}=\{f_{2}^{\text{GAG}},f_{2}^{\text{G\&M}}\}$. Following [lacasa2025certification], we consider a split to be adequate if (i) finding the same data samples appearing in both the train and test sets has vanishingly small probability, (ii) the feature vector distributions of both sets are similar, and (iii) the test set data is adequately close from the training set.
Criterion (ii) is checked by performing 2-sample hypothesis tests (2-sample Kolmogorov-Smirnov (KS) [scholz1987k], 2-sample Anderson-Darling (AD) [hodges1958significance] and 2-sample chi-2 tests) on marginal training and test distributions for each input variable. Criterion (iii) is implemented in the so-called tesselation and proximity method [lacasa2025certification], by detecting test set points that are either excessively close (statistically speaking) to the training set (source of p-hacking), or statistically too far (isolated points with poor generalization).
@@
Figure 17: Training vs test marginal distributions for the output variable NN.
In Figure 15 we plot the marginal training and test set distributions for each numerical input variable, whereas in Figure 16 we plot the ones for the categorical variables. Although variables in Figure 16 are actually numerical, for the sake of this analysis they are treated as categorical as they only take a small set of possible values. Therefore, they cannot be considered continuous and a different test is applied ($\chi^{2}$). Likewise, Figure 17 presents the marginal training and test distributions for the output variable NN. Visually, the training and test distributions are similar, although small differences exist. To quantify the significance of these, we have run several hypothesis tests, where the null hypothesis H0 is that, for each variable, each pair of populations are drawn from the same distribution. Since each hypothesis test has its own null hypothesis, no correction for multiple testing is required, and we adopt the standard choice for rejecting the null. In Table 7 we present the p-values of these tests, where the null hypothesis cannot be rejected for any variable except for $n$ and $t_{flight}$. Note nonetheless that, as observed in Figure 16, the range values for these two variables is similar in both sets, indicating that at least there are no isolated values in $n$ or $t_{flight}$ in the test set.
@@ 4 Conclusions
Looking ahead, while the current study is limited to wing components due to data availability, the methodology is extensible to other materials and aircraft structures, such as the fuselage and stabilizers. Future work should also focus on broader structural coverage, improved stress modeling, modifications to the pipeline aimed at promoting conservativeness, (e.g., using modified loss functions to penalize non-conservative predictions), and further integration of uncertainty quantification conditioned on input space [lacasa2025certification].
@@
Acknowledgments – The authors acknowledge funding from project TIFON (PLEC2023-010251) funded by MCIN/AEI/10.13039/501100011033, Spain. $L$ acknowledges partial support from projects MISLAND (PID2020-114324GB-C22), project CSxAI (PID2024-157526NB-I00) funded by MICIU/AEI/10.13039/501100011033/FEDER, UE, project MdM Seal of Excellence (CEX2021-001164-M) funded by MICIU/AEI/10.13039/501100011033 and from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL). Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer.
@@ B.5 $f_{2}^{\text{G\&M}}$: input, output variables and model specs
In this case we only have 7 input variables: $k_{t},\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4},n,t_{flight}$. The rationale of each variable is similar to that for $f_{2}^{\text{GAG}}$; $\langle 1g\rangle_{\text{flight}}$ shows the stress level around which the oscillations that cause this damage take place. The magnitude of these oscillations is captured by $\langle{\Delta\text{vgust}}\rangle,\langle{\Delta\text{vman}}\rangle,\langle{\Delta\text{turn}}\rangle$ and $n$, $t_{flight}$ have the same statistical impact. Besides, longer flights typically imply more gusts encountered and maneuvers executed (not necessarily larger) and consequently greater damage. The output variable is the per-flight damage $\overline{\mathscr{D}_{\text{G\&M}}}$, such that the cumulated damage is simply $\mathbb{D}_{\text{G\&M}}=n\cdot\overline{\mathscr{D}_{\text{G\&M}}}$. Once again, variables were normalized with MinMax Scaler and logarithm was applied to the damage. The MLP specs and hyperparameters are specified in Table 14. Hyperparameter tuning was performed manually with the validation set to ensure generalization.
@@
Appendix C Statistical validation of moment closure and finite sample deviations
As stated in Section 2, our method predicates on approximating the individual damages in each flight –the latent random variables $\mathscr{D}_{\text{GAG}}$ and $\mathscr{D}_{\text{G\&M}}$ – by their sample mean. In this appendix, we give insight into the validity of the implicit assumptions behind this procedure with an analysis based on the GAG damage.
There are mainly two approximations:
  • 1.
    First, we identify each realization of the GAG damage random variable $\mathscr{D}_{\text{GAG}}$ with the observed per-flight average $\overline{\mathscr{D}_{\text{GAG}}}$, i.e, we approximate a random variable with its (estimated) first moment. This is sometimes called a mean-field approximation or a first-moment closure, and it is a valid approximation when higher moments of the distribution (e.g. fluctuations of the random variable around its mean value) are relatively small (in particular, much smaller than the mean). Such comparison can be made by computing the so-called squared coefficient of variation, $\text{CV}^{2}=\frac{\sigma^{2}}{\overline{\mathscr{D}_{\text{GAG}}}^{2}}$, where $\sigma^{2}$ and $\overline{\mathscr{D}_{\text{GAG}}}$ are the sample variance and sample mean of the GAG damage per-flight, respectively. First moment closures usually require $\text{CV}^{2}\ll 1$. In Table 15 we provide an illustrative table with results for three missions for a given PSE and $k_{t}$. As observed, identifying each realization of the per-flight random variable $\mathscr{D}_{\text{GAG}}$ with its sample average is not fully justified yet, $\text{CV}^{2}\leq 1$, not $\text{CV}^{2}\ll 1$, suggesting non-negligible fluctuations. However, as specified in section 2 the damage of one particular flight is not the variable of interest. On the contrary, the important variable is the cumulated damage after $n$ flights, $\mathbb{D}_{\text{GAG}}=\sum_{i=1}^{n}\mathscr{D}_{\text{GAG}}$. Accordingly, the first moment closure needs to hold with respect to this cumulated random variable $\mathbb{D}_{\text{GAG}}$. Now, for large $n$ (when expected values coincide with sample averages), the expected value of the cumulated GAG damage $\mathbb{E}(\mathbb{D}_{\text{GAG}})\equiv\overline{\mathbb{D}_{\text{GAG}}}$, which for ease of notation we label it $\mathbb{D}_{\text{GAG}}$, reads $\mathbb{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$, since
    $\mathbb{E}(\mathbb{D}_{\text{GAG}})=\mathbb{E}(\sum_{i=1}^{n}\mathscr{D}_{\text{GAG}})=\sum_{i=1}^{n}\mathbb{E}(\mathscr{D}_{\text{GAG}})=n\overline{\mathscr{D}_{\text{GAG}}}.$ (7)
    At the same time, as we are assuming flight independence we have
    $\text{Var}(\mathbb{D}_{\text{GAG}})=\text{Var}(\sum_{i=1}^{n}\mathscr{D}_{\text{GAG}})=\sum_{i=1}^{n}\text{Var}(\mathscr{D}_{\text{GAG}})=n\cdot\sigma^{2}.$ (8)
    Accordingly, assessment of the first moment closure approximation via the squared coefficient of variation reads now
    $\text{CV}^{2}_{\mathbb{D}_{\text{GAG}}}=\frac{\text{Var}(\mathbb{D}_{\text{GAG}})}{\mathbb{E}(\mathbb{D}_{\text{GAG}})^{2}}=\frac{n\cdot\sigma^{2}}{n^{2}\cdot\overline{\mathscr{D}_{\text{GAG}}}^{2}}=\frac{\text{CV}^{2}}{n}$ (9)
    In short, the first moment closure is a substantially improved approximation for large $n$ (see Table 15 for three different missions), and overall the procedure is thus sound and justified.
    $n$ $\text{CV}^{2}$ $\text{CV}^{2}_{\mathbb{D}_{\text{GAG}}}$
    $\approx 800$ 1 0.001
    $\approx 1300$ 0.3 0.0002
    $\approx 13000$ 0.09 0.000007
    Table 15: Squared coefficients of variation for $\mathscr{D}_{\text{GAG}}$ and $\mathbb{D}_{\text{GAG}}$, for three different PSEs illustrating very different number of flights $n$.
  • 2.
    Second, approximating the (true, hidden) expected value of a random variable by its sample mean is a classical estimation method, which holds by virtue of the law of large numbers which assures almost-surely convergence of the sample mean to the true mean for large sample sizes. For illustration, we assess how the (true, hidden) average per-flight GAG damage deviates from the (estimated) sample average per-flight GAG damage $\overline{\mathscr{D}_{\text{GAG}}}$ using non-parametric bootstrapping and building the 95% confidence interval of the bootstrapped mean GAG Damage, as this gives us an uncertainty interval where the true mean will lies with high probability. In Figure 18 we represent, for a specific PSE, $k_{t}$ and mission ($n\approx 1300$ flights), the distribution of bootstrapped means alongside its 2.5 and 97.5 percentiles as well as the bootstrap average. We also locate the original sample mean for reference. Observe that the sample mean of the GAG damage, $\overline{\mathscr{D}_{\text{GAG}}}$ and the average of the bootstrapped means $\mu$ virtually coincide, $\overline{\mathscr{D}_{\text{GAG}}}\approx\mu=6.19\times 10^{-5}$, and the bootstrap $CI95=[6.10\times 10^{-5},6.29\times 10^{-5}]$ is a very narrow one. This validates our procedure of identifying the true mean with the sample mean. Results hold similarly for other cases.
Figure 18: Distribution of bootstrapped means associated to the GAG damage $\mathscr{D}_{\text{GAG}}$, for a specific PSE and mission. The 2.5 and 97.5 percentiles –conforming the bootstrap CI95 of average per-flight GAG damage are highlighted, alongside the average of the bootstrapped means $\mu$ and the sample mean $\overline{\mathscr{D}_{\text{GAG}}}$.
This procedure critically depends on the number of flights $n$ being moderately large, something that in this work systematically holds as we have $n\in[800,13000]$. Now, it is always desirable to establish what would be the minimum number of flights $n_{\text{min}}$ such that our method is acceptable. This is akin to ask for what sample size (i) the first moment closure is no longer a valid one, or (ii) when estimating the true mean with the sample mean is not acceptable. The actual question is difficult to answer in general, as it requires to define what we mean by ‘acceptable’. Here we provide two different roadmaps to estimate the minimum $n_{min}$ to meet a certain statistical guarantee, namely, that with a certain significance level $\alpha$ and confidence level $p=1-\alpha$, the mismatch between the sample mean of the per-flight GAG damage $\overline{\mathscr{D}_{\text{GAG}}}$ and the true mean is smaller than a certain tolerance $\epsilon$ (measured in relative error units).
In general terms, invoking the central limit theorem (CLT, which holds as high order moments of $\mathscr{D}_{\text{GAG}}$ do not diverge, see above), such difference drops to zero as fast as $1/\sqrt{(}N)$. In the event that $\mathscr{D}_{\text{GAG}}$ is Gaussian, simple probability arguments provide a closed expression for $n_{min}$:
$n_{min}=\left(\frac{Z(\alpha/2)\sigma}{\epsilon\overline{\mathscr{D}_{\text{GAG}}}}\right)^{2},$ (10)
where $\alpha$ is our desired significance level, $Z(\alpha/2)$ can be obtained from the values from a standard normal distribution, and sample standard deviation $\sigma$ and the sample mean $\overline{\mathscr{D}_{\text{GAG}}}$ can be computed from the population dataset. When the marginal distribution deviates from a pure Gaussian, the expression above is only an approximation, which nonetheless becomes a tight one as soon as $n$ is moderately large (our case) and if the underlying distribution does not have strong fluctuations around the mean (i.e. long tails), which, again, is the case here.
For illustration, for the concrete mission and PSE discussed above with $n\approx 1300$, and fixing $\epsilon=0.02$ and $\alpha=0.05$ (i.e. allowing for a maximum 2% relative error with a 95% confidence), we find $n_{min}$ = 883, which is substantially smaller than the actual number of flights. Similarly, we can estimate the maximum error associated with the exact $n$ used. In this case, for a 95% confidence level, the maximum error is $1.64\%$.
Finally, we can also define an heuristic which makes use of nonparametric bootstrapping (and therefore does not assume any shape for the GAG Damage distribution). The minimal acceptable size of the dataset can be found by (i) establishing an acceptable interval for the sample mean in the original dataset (e.g. the CI95 of the bootstrapped mean), then (ii) sequentially downsampling the original dataset, estimating the new sample mean in each case, and (iii) detecting when the sample mean is no longer within the established interval.
In practice, the procedure is implemented as follows: first, we define a new data size $n_{new}<n$. Then, we perform independent resamplings (here we performed 2000, with replacement) of the original dataset of size $n_{new}$, and compute the percentage of cases where the new sample mean of the resampled datasets falls within the originally established interval.
These ‘acceptable intervals’ were built with two different approaches, each providing different and complementary information: i) the interval is defined by fixing a priori a relative error around the average of the bootstrapped means $\mu$ (see e.g. Figure 18) so as to build the interval $\mu-\epsilon\mu,\mu+\epsilon\mu]$, and ii) the interval is defined as the 95% confidence interval of the bootstrapped mean.
For illustration, using the same mission and PSE with $n\approx 1300$, in Figure 19 we show the percentage of subsets falling within each interval respectively, as a function of $n_{new}.$ Observe that for a sample size of 883 (flagged as $n_{\text{min}}$ by the analytical approach above), in about $95\%$ of the resamples the sample average error is equal or lower than the established error, (2%), whereas $87.5\%$ of the samples fall in the CI95 interval.
Figure 19: (Left panel) Percentage of resamples with $n_{\text{new}}<n$ data where the sample mean is within the acceptable interval defined initially as $\mu-\epsilon\mu,\mu+\epsilon\mu]$, as a function of the datasize sample $n_{\text{new}}$. (Right panel) Similar to the left panel, but here the acceptable interval is defined as the bootstrap mean CI95.